OpenStudy (anonymous):

Simplify to only x: x+8(sqtx+8)

jhonyy9 (jhonyy9):

sqrtx+8 there 8 is inside radical too ?

OpenStudy (anonymous):

\[x+8\sqrt{x+8}\]

jhonyy9 (jhonyy9):

than yes so x+8(sqrt(x+8))=0 8(sqrt(x+8)= -x squared both side 64(x+8)=x*x 64x+512=x2 x2-64x-512=0

jhonyy9 (jhonyy9):

x2 is x squared